OpenStudy (anonymous):

what is the solution to this system of linear equations? y=-8x-6 y=-2x+6

OpenStudy (anonymous):

x=-2 y=6

OpenStudy (anonymous):

thats not right

OpenStudy (anonymous):

sorry y=10

OpenStudy (anonymous):

\[\{x=-2,y=10\} \]

OpenStudy (anonymous):

what about y=6x+52 and y=3x+28

OpenStudy (anonymous):

\[\{x=-8,y=4\} \]

OpenStudy (anonymous):

y=-6x+52, y=2x+17

OpenStudy (anonymous):

\[\left\{x=\frac{35}{8},y=\frac{103}{4}\right\} \]